1.the following information are given XYZ manufacturing company:-
- fixed costs=$35550
-variable cost per unit=50(x)
-the production volume(capacity) is 10000 units
- R(x)=80x
Required:-
based on the above information, compute the break-even point
A, in units
B, as percent of capacity
c, in dollar
1
Expert's answer
2020-07-20T18:07:23-0400

A) "35,550+50x=80x"

"x=1185"

B) "1185/10,000*100=11.85\%"

C) "1185*80=94,800" $